OpenStudy (anonymous):

Find x : e^x e^(x+1) = 1 will give medals!!

OpenStudy (anonymous):

(e^x)(e^(x+1)) = e^(2x+1) [Law of Exponents] So e^(2x+1) = 1 The trick is to write 1 as e^0. so you have e^(2x+1) = e^0 so equate the exponents (since both bases are e) therefore, 2x + 1 = 0 x = -1/2 @QueenBee232